A homogeneous good industry is composed of 3 firms. You are given the following information on output, price and marginal cost of each firm: q, 200 q2-500 93100 p- 50 41.7 c2 29.2 c3 45.8 Remember that for each firm where α, is the market share of firm i and η is the price elasticity of demand. a)Calculate the 2-firm concentration ratio. b)Calculate the Herfindahl index c) Calculate the number equivalent. d) Calculate the Lerner index of ach firm. e Calculate the price elasticity of demand. a) The bicycle industry consists of seven firms. Firms 1, 2, 3, 4 each has 10% market share, and firms 5,6,7 each has 20% market share. Using the concentration measures, answer the following questions (i) Calculate 4-firm concentration ratio for this industry (ii) Calculate the Herfindahl index (IHH) for this industry (iii) Now, suppose that firms 1 and 2 merge, so that the new firm will have a market share of 20%
Answer #1

Answering only first question is mandatory by HomeworkLib policy

Q1) Q = q1 + q2 + q3

= 200+500+100

= 800

So share of 1st firm S1 = 200/800 = 1/4 = 25%

S2 = 500/800 = 5/8 = 62.5%

S3 = 100/800 = 1/8 = 12.5%

So C2 = S2 + S1 = 62.5 + 25

C2 = 87.5%

B) HHI = \sumsi2

= .252+ .6252 + .1252

= .46875

C) number equivalent = 1/ HHI

= 1/.46875

= 2.13333

D) Lerner index =(P-MC)/P

For first firm L1 = (50-41.7)/50 = .166

L2= .416

L3 = .084

E) (50-41.7)/50 = .25/n

Elasticity of demand ,n = .25/.166

n= 1.506